OpenStudy (anonymous):

A delivery service charges a base price for an overnight delivery of a package plus an extra charge for each pound the package weighs. A customer is billed $22.85 for shipping a 3-pound package and $40 for shipping a 10-pound package. 1.) Write an equation that gives the total cost of shipping a package as a function of the weight of the package. 2.)Find the cost of shipping a 15-pound package

OpenStudy (ash2326):

@Jon1122 Let the base price be x and price per pound be y, can you frame an equation for this info. A customer is billed $22.85 for shipping a 3-pound package

OpenStudy (anonymous):

So it would be: Price= Pound(y) + x

OpenStudy (anonymous):

X=baseprice

OpenStudy (ash2326):

Good, yes, that's the basic equation. We need to find the value of x and y Now for the first shipment, customer pays 22.85 for 3 pound, can you plugin the values here?

OpenStudy (anonymous):

I really have no idea, sorry. If I'm wasting your time you should move onto something else.

OpenStudy (ash2326):

no problem dude, I'm here to help.

OpenStudy (anonymous):

Do you already know the answer?

OpenStudy (ash2326):

We have the base equation as Total Price= Base Price + Price/pound Here we have \[Total Price= x+ y \times P\] for the first case P=no. of pounds =3 Total Price= 22.85 so we have \[22.85=x+3\times P\]

OpenStudy (ash2326):

do you follow? No I don't know the answer already

OpenStudy (anonymous):

Ok so what's next?

OpenStudy (ash2326):

now you use the second info $40 for shipping a 10-pound package. can you plugin the values like I did before to make an equation?

OpenStudy (anonymous):

40 = 3(x) X p

OpenStudy (ash2326):

check again

OpenStudy (anonymous):

40 = 3+x times p?

OpenStudy (ash2326):

40=x+3y

OpenStudy (ash2326):

do you follow?

OpenStudy (anonymous):

Yeah

OpenStudy (ash2326):

okay, do you know how to solve equations with two variables?

OpenStudy (anonymous):

Yes

OpenStudy (ash2326):

could you try to find the value of x and y?

OpenStudy (anonymous):

No idea....

OpenStudy (anonymous):

3 does no go evenly into 40

OpenStudy (anonymous):

Ugh I'm so confused :(